1.      Two examples are randomly selected from two groups of students who have been taught using different teaching methods. An examination is given and the results are shown below:

Group 1                               Group 2

N1 = 8                                  n2 = 10

Mean of Grp 1 =85         Mean of Grp 2 =87

Variance of Grp 1 =46    Variance of Grp 2 = 36.

               Using a .05 level of significance, can we conclude that the two different teaching methods are equally effective?


1
Expert's answer
2021-03-18T14:01:05-0400

"H_0 : μ_1 = μ_2 \\

H_1 : μ_1 ≠ μ_2 \\

n_1 = 8 \\

\bar{X_1} = 85 \\

s^2_1 = 46 \\

n_2 = 10 \\

\bar{X_2} = 87 \\

s^2_2=36"

Because both samples are small (< 30), we use the t test statistic. Before implementing the formula, we first check whether the assumption of equality of population variances is reasonable. The ratio of the sample variances, "\frac{s_1^2}{s_2^2} = \frac{46}{36} = 1.27" , which falls between 0.5 and 2, suggesting that the assumption of equality of population variances is reasonable. The appropriate test statistic is:

"t = \frac{\bar{X_1} – \bar{X_2}}{S_p \sqrt{ \frac{1}{n_1} + \frac{1}{n_2}}}"

This is a two-tailed test, using a t statistic and a 5% level of significance. The appropriate critical value can be found in the t Table. In order to determine the critical value of t we need degrees of freedom, df, defined as df=n1+n2-2 = 8+10-2=16. The critical regions for a two-tailed test with df=16 and α=0.05 are t ≥ 1.745 and t ≤ -1.745 and the decision rule is: Reject H0 if t ≥ 1.745 or t ≤ -1.745.

We now substitute the sample data into the formula for the test statistic. Before substituting, we will first compute Sp, the pooled estimate of the common standard deviation.

"S_p = \sqrt{ \frac{(n_1-1)s^2_1+(n_2-1)s^2_2}{n_1+n_2-2} } \\

S_p = \sqrt{ \frac{(8-1)46+(10-1)36}{8+10-2} } \\

S_p = 6.35 \\

t = \frac{85 – 87}{6.35 \sqrt{ \frac{1}{8} + \frac{1}{10}}} = -0.66"

Since -0.66 > -1.745, we cannot reject the null hypothesis.

Using a 0.05 level of significance, we can conclude that the two different teaching methods are equally effective.